Download HW 15 (Excel download) and perform the necessary MRP calculations. Then enter selected values below.

What is the Parent's Net Requirements for Period 5?

What is the Parent's Net Requirements for Period 8?

What is the Parent's Planned Order Receipts for Period 6?

What is Part 1's Net Requirements for Period 8?

What is the Part 1's Net Requirements for Period 5?

What is the Part 1's Planned Order Release for Period 5?

What is Part 2's Gross Requirements for Period 7?

What is Part 2's Projected On-Hand Inventory for Period 2?

What is Part 2's Net Requirements for Period 6?

What is Part 2's Planned Order Receipt for Period 6?

Management Instructions
Beginning Inventory Lot Size Lead Time Quantity per Parent Safety Stock
α μ σ Amount
Parent 403 12 1 93.0% 5,300 40
Part 1 64 12 2 9 50.0% 1,900 30
Part 2 -    200 1 7 80.0% 675 5
Material Requirements Plan
Period 1 2 3 4 5 6 7 8
Demand -    -    -    134 143 60 -    49
Parent
Gross Requirements
Scheduled Receipts 14
Projected On-Hand
(at beginning of period)
Net Requirements
Planned Order Receipts
Planned Order Releases
Part 1
Gross Requirements
Scheduled Receipts
Projected On-Hand
(at beginning of period)
Net Requirements
Planned Order Receipts
Planned Order Releases
Part 2
Gross Requirements
Scheduled Receipts 85 85
Projected On-Hand
(at beginning of period)
Net Requirements
Planned Order Receipts
Planned Order Releases
Answer #1

EXCEL FORMULAS:

Management Instructions
Beginning Inventory Lot Size Lead Time Quantity per Parent Safety Stock
α μ σ Amount
Parent 403 12 1 0.93 5300 40 =NORMSINV(G4)*I4*SQRT(E4)
Part 1 64 12 2 9 0.5 1900 30 =NORMSINV(G5)*I5*SQRT(E5)
Part 2 200 1 7 0.8 675 5 =NORMSINV(G6)*I6*SQRT(E6)
Materials Requirements Plan
Period 1 2 3 4 5 6 7 8
Demand 134 143 60 49
Parent
Gross Requirements =C10 =D10 =E10 =F10 =G10 =H10 =I10 =J10
Scheduled Receipts 14
Projected On-Hand (at beginning of period) =C4+C14+C17 =C15-C13+D14+D17 =D15-D13+E14+E17 =E15-E13+F14+F17 =F15-F13+G14+G17 =G15-G13+H14+H17 =H15-H13+I14+I17 =I15-I13+J14+J17
Net Requirements =MAX(0,C13-C15) =MAX(0,D13-(C15-C13+D14)) =MAX(0,E13-(D15-D13+E14)) =MAX(0,F13-(E15-E13+F14)) =MAX(0,G13-(F15-F13+G14)) =MAX(0,H13-(G15-G13+H14)) =MAX(0,I13-(H15-H13+I14)) =MAX(0,J13-(I15-I13+J14))
Planned Order Receipts =IFERROR(1*OFFSET(D18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(E18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(F18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(G18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(H18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(I18,0,-1*$E$4,1,1),0) =IFERROR(1*OFFSET(J18,0,-1*$E$4,1,1),0)
Planned Order Releases 0 0 0 0 0 0 0 0
Part 1
Gross Requirements =C18*$F$5 =D18*$F$5 =E18*$F$5 =F18*$F$5 =G18*$F$5 =H18*$F$5 =I18*$F$5 =J18*$F$5
Scheduled Receipts
Projected On-Hand (at beginning of period) =C5 =C23-C21+D22+D25 =D23-D21+E22+E25 =E23-E21+F22+F25 =F23-F21+G22+G25 =G23-G21+H22+H25 =H23-H21+I22+I25 =I23-I21+J22+J25
Net Requirements =MAX(0,C21-C23) =MAX(0,D21-(C23-C21+D22)) =MAX(0,E21-(D23-D21+E22)) =MAX(0,F21-(E23-E21+F22)) =MAX(0,G21-(F23-F21+G22)) =MAX(0,H21-(G23-G21+H22)) =MAX(0,I21-(H23-H21+I22)) =MAX(0,J21-(I23-I21+J22))
Planned Order Receipts =IFERROR(1*OFFSET(D26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(E26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(F26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(G26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(H26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(I26,0,-1*$E$5,1,1),0) =IFERROR(1*OFFSET(J26,0,-1*$E$5,1,1),0)
Planned Order Releases =ROUNDUP((OFFSET(C24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(D24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(E24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(F24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(G24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(H24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(I24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5 =ROUNDUP((OFFSET(J24,0,$E$5,1,1)+$J$5)/$D$5,0)*$D$5
Part 2
Gross Requirements =C26*$F$6 =D26*$F$6 =E26*$F$6 =F26*$F$6 =G26*$F$6 =H26*$F$6 =I26*$F$6 =J26*$F$6
Scheduled Receipts 85 85
Projected On-Hand (at beginning of period) =C6 =C31-C29+D30+D33 =D31-D29+E30+E33 =E31-E29+F30+F33 =F31-F29+G30+G33 =G31-G29+H30+H33 =H31-H29+I30+I33 =I31-I29+J30+J33
Net Requirements =MAX(0,C29-C31) =MAX(0,D29-(C31-C29+D30)) =MAX(0,E29-(D31-D29+E30)) =MAX(0,F29-(E31-E29+F30)) =MAX(0,G29-(F31-F29+G30)) =MAX(0,H29-(G31-G29+H30)) =MAX(0,I29-(H31-H29+I30)) =MAX(0,J29-(I31-I29+J30))
Planned Order Receipts =IFERROR(1*OFFSET(D34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(E34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(F34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(G34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(H34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(I34,0,-1*$E$6,1,1),0) =IFERROR(1*OFFSET(J34,0,-1*$E$6,1,1),0)
Planned Order Releases =ROUNDUP((OFFSET(C32,0,$E$6,1,1)+$J$6)/$D$6,0)*$D$6 =ROUNDUP((OFFSET(D32,0,$E$6,1,1)+$J$6)/$D$6,0)*$D$6 =ROUNDUP((OFFSET(E32,0,$E$6,1,1)+$J$6)/$D$6,0)*$D$6 =ROUNDUP((OFFSET(F32,0,$E$6,1,1)+$J$6)/$D$6,0)*$D$6 =ROUNDUP((OFFSET(G32,0,$E$6,1,1)+$J$6)/$D$6,0)*$D$6 0 0 0